# AugGen: Augmenting Task-Based Learning in Professional Creative Software with LLM-Generated Scaffolded UIs

### v1

@@ 5.2. Study 2: Expert Users
For the study with experts, we kept RQ1–RQ3 for a comparative analysis across beginner and expert users. In addition, two further research questions were formulated to explore the impact on skill acquisition and method generalization.
@@
  • RQ4: How can the scaffolded UI potentially shape long-term skill development? This question invites expert perspectives on whether the scaffolded interface can help refine skills, adapt tools to new tasks, and strengthen concept understanding for solving new problems.
@@
  • RQ5: What is the potential impact of applying the scaffolded UI generation method to other software? This question asks for insights from experts familiar with additional professional creative software beyond Blender on the benefits and challenges of using scaffolded interface in different platforms.
@@
Experiment design. Each participant was assigned to one of two task conditions. Task 1 involved performing UV unwrapping on a 3D monkey head (Figure 8), an advanced texturing task, while Task 2 required constructing a human walk cycle that reflected personality, an open-ended animation task. We provided the 3D model for Task 1 and the rigged human character for Task 2 for experiment consistency. Task assignment was based on participant prior experience: only those with 3D model texturing experience were assigned to Task 1, and only experienced animators were assigned to Task 2. This allowed us to gather feedback grounded in a participant’s prior expertise to enable comparisons between their experience with the Blender interface and the scaffolded interface.
@@
Participant recruitment. We recruited eight experts (3 female, 5 male; mean age = 28.87, SD = 3.66) via Prolific (prolific), requiring at least one year of Blender experience and preferring additional proficiency in tools like Maya or 3ds Max. All other criteria for gender, age, language, and Prolific history matched those of Study 1. Participants (S2-P1 to S2-P8) were each assigned to either Task 1 or Task 2 and completed their task using the scaffolded interface, with four experts per task. Detailed demographics appear in Appendix C.
@@
Measurements. To evaluate RQs 1, 2, and 3, we used the same questionnaires as Study 1 and adapted the items to preference-based questions: each item was rated on a 7-point Likert scale (1 = strongly prefer our interface, 7 = strongly prefer the baseline interface). Additionally, we collected qualitative feedback by asking participants to elaborate on their experience with our interface regarding task performance, concept learning, and interface experience. For RQs 4 and 5, we gathered user qualitative data regarding skill development and using the scaffolded interface in other software tools. Refer to Appendix B.4 for the open-ended questions for qualitative data collection.
@@
Procedure. We invited eligible participants to our study via Zoom. The pre-study session involving the consent and pre-study questionnaire had the same procedure as Study 1. Following this, we offered a brief introduction to the user’s task ($\sim$5 minutes). After participants verbally confirmed that they were clear about the task requirements, they used the scaffolded interface to work on the task for 15 minutes through remote desktop control. We recorded the desktop screen with participant consent and took notes on their actions. After the task, participants completed the post-study questionnaires and answered the open-ended questions ($\sim$10 minutes). The study concluded with a payment of 10 USD for completing the 30-minute study.
@@
Data analysis. The qualitative data was analyzed using thematic analysis (braun2006using) with a deductive strategy. The authors reviewed the participant responses, coded the data to extract relevant information centered around the research questions, refined the codes, and categorized themes.
@@ 5.3. Study Design Remarks
Since our study with beginners aimed to explore their learning experience using scaffolded UIs in a complex creative tool rather than making strong causal claims, not a large sample size was the main priority. We used a between-subjects design to avoid cross-condition contamination during learning and skill acquisition. In the study with experts, we prioritized qualitative depth over large-scale inference, and the eight users frequently interacted with Blender and were able to articulate pros and cons of our scaffolded interface clearly. Also, we observed signs of saturation during thematic analysis of the qualitative data.
@@ 6.1.1. Questionnaire Responses
Figure 9 shows the participant reported responses to the questionnaires. Asterisks in the figure indicate statistical significance derived from post-hoc pairwise comparisons between Ours (scaffolded) and Baseline (default Blender) interfaces for each task and measure. The complete two-way ANOVA results are reported in Appendix D. From the results, we see that our scaffolded interfaces have significantly improved task performance, concept learning, and user experience compared to the default Blender interface.
@@
(c) Concept learning.
@@
Figure 9. Study 1 participant responses using Ours (scaffolded) and Baseline (default Blender) interfaces for Tasks 1 and 2 (lower is better). Asterisks denote significant differences in post-hoc pairwise comparisons between interfaces within each measure (*: $p<.05$, **: $p<.01$, and ***: $p<.001$).:
@@ 6.1.2. Task Completion Time
All participants successfully completed their tasks and generally spent less time using our scaffolded interface compared to the default Blender interface: $\bar{t}_{Ours(Task1)}=12.69\pm 1.49$ mins vs. $\bar{t}_{Baseline(Task1)}=15.08\pm 2.34$ mins; $\bar{t}_{Ours(Task2)}=20.39\pm 2.55$ mins vs. $\bar{t}_{Baseline(Task2)}=24.95\pm 3.46$ mins. Significant main effects of Interface ($F_{(1,30)}=14.78,p<.001,\eta^{2}_{p}=.33$) and Task ($F_{(1,30)}=94.71,p<.001,\eta^{2}_{p}=.76$) were observed. The interaction effect was not significant. Post-hoc analysis revealed no significant difference in completion time in Task 1, but users with our interface were significantly faster in Task 2 ($p<.05$), demonstrating the efficiency advantage of our scaffolded interface under higher task demands.
@@ 6.1.3. User Behavior
We observed that participants using our scaffolded interface (experiment group) progressively reduced their reliance on tutorial videos and increasingly navigated via the interface’s inherent workflow structure, echoed by participant self-reported improved workflow clarity, tool relevance, and reduced external help requirement. In contrast, participants using the default Blender interface (control group) consistently referred to the videos throughout the tasks. Consequently, the experiment group generally completed tasks faster than the control group. Additionally, when some tools were not mentioned in the tutorials but presented in our scaffolded interfaces, the experiment group explored them during or after their tasks, as echoed by participant self-reported tool and concept exploration. Conversely, the control group rarely explored beyond the tools covered in tutorials.
@@ 6.2.1. Questionnaire Responses
Figure 10 shows the participant reported preference scores for Ours (scaffolded) or Baseline (default Blender) interfaces on the measures of task performance, concept learning, and user interface experience. They had a clear preference for our interface across all dimensions regarding task performance. For concept learning, participants generally perceived our interface as more supportive for understanding concepts, relating tools to concepts and workflows, and motivating learning. However, they had mixed opinions on concept exploration (4 favored our interface, 1 was neutral, 3 favored the baseline). For the overall interface experience, consistent preference for our interface was observed except for tool exploration (5 participants preferred our interface, 3 favored the baseline).
Figure 10. Study 2 participant preferences for Ours (scaffolded) or Baseline (default Blender) interfaces regarding task performance, concept learning, and user interface experience.:
@@ 6.2.2. Qualitative Results
Tradeoffs of supporting task performance and allowing workflow flexibility (RQ1). Participants generally appreciated task performance support by surfacing task-relevant tools (S2-P1, S2-P2, S2-P7, S2-P8). The scaffolded interface led to higher task efficiency and user focus, as S2-P2 said that, “I actually finished the UV task faster than expected.” However, participants also expressed concerns that the system’s assumptions sometimes misaligned with their preferred workflows. Some found the interface limiting when it failed to surface user-expected tools (S2-P4). Others perceived the guided structure forced a specific process, which may reduce workflow flexibility (S2-P6). These findings indicate the need for more user customization to support expert users who prefer flexible or iterative workflows.
@@
Gains of simplicity and focus vs. loss of familiarity (RQ3). Participants responded positively to the overall experience since the scaffolded interface was “clean, not overwhelming” (S2-P1), “task-centered” (S2-P4), “lightweight and approachable” (S2-P7), and with “limited distractions” (S2-P8). S2-P2 also noted that, “I can imagine a less fighting the UI if I were a newbie,” indicating reduced cognitive friction during software navigation for beginners. Nevertheless, participants commented that the scaffolded interface may conflict with their familiarity with the default software layout. S2-P3 remarked that they “missed the old familiarity,” pointing to the tension when introducing novel interfaces.
Scaffolding’s impact on skill development (RQ4). Participants found that the scaffolded interface provided a supportive foundation for building practical skills, particularly for newcomers. For example, S2-P5 called it a “great stepping stone from beginner to intermediate,” and S2-P6 noted the interface made it “easier to build confidence in doing simple stuff.” These responses show that experts see our interface as helping beginner users build confidence and gain skills while learning to use professional creative software.
Using scaffolded interfaces in broader contexts (RQ5). Participants with experience in other professional creative software recognized the underlying principles of scaffolded interface design as being transferable. For instance, S2-P8 stated that, “I think the knowledge about motion, timing, bone constraints translates to using other tools.” Building on such transferability, participants suggested potential applications and impacts of using scaffolded interfaces in other software. S2-P6 said that, “3ds Max can benefit from this. The modifier stack alone is complex enough. If tools were grouped by task and revealed based on what you’re actually doing, it’d cut down on confusion.” Participants also anticipated the benefits, e.g., “it could help new users get productive faster, and even experienced users might find it useful to stay focused” (S2-P6). These results suggest user enthusiasm for applying the scaffolded interface generation method more broadly to reduce interface complexity and enhance workflow efficiency.
@@ 6.2.3. User Behavior
We observed that expert users approached tasks flexibly, sometimes bypassing the step-by-step workflow stages in the scaffolded interface. For instance, S2-P6 switched between marking seams (stage 1) and unwrapping & editing (stage 2). They started with the “Select Edge Loops” tool in stage 1 to mark seams, used “Unwrap” in stage 2 to check the UV map, then went back to use “Edge Select” in stage 1 to refine the seams and repeated the process. This reflects their preference for multitasking and iterative workflows. Furthermore, when needed tools were not presented at their current complexity level, several participants (S2-P2, S2-P5, S2-P6, S2-P8) sometimes did not look for these tools by advancing complexity levels. Instead, they accessed tools via keyboard shortcuts or the default Blender interface—the operations they were familiar with. This aligns with their feedback that the scaffolded interface was less supportive for tool exploration compared to the experience of beginners.
@@ 7. Discussion
Based on the user study findings, here we discuss the opportunities in AugGen, its limitations, and future directions.
7.1. Opportunities in AugGen
Task-aware adaptation to support software learners and educators. In our study, beginner users reported benefits in both task performance and concept learning with our scaffolded interface, showing its promise for instructional use. To better support learning for software learners, the interface can adapt its scaffolding strategy based on task complexity. For simple tasks, the interface can serve as a guided playground where future design could expand the available tools slightly beyond the current step to promote curiosity-driven exploration. For complex tasks, the focus shifts to practice and learning, so the interface can offer a smaller, essential toolset with clear links between domain concepts and software operations. It could also include best-practice tips or visual aids like video clips to help ensure key concepts and operations are grasped before moving on.
@@ 7.1. Opportunities in AugGen
Scaffolded interface driven by productivity and creativity. Our study with experts revealed that they focused on boosting productivity through personalized workflows and creativity by flexible exploration. Therefore, scaffolded interfaces targeting at expert end users should evolve beyond structured guidance towards systems driven by these needs. To support productivity, a scaffolded interface could learn a user’s unique shortcuts, tool sequences, and niche commands, and optimize workflows by rearranging palettes for top tools or suggesting more efficient alternatives. Implementing this requires robust user activity logging, potentially through screen monitoring by in-application AI copilots (cunningham2025copilotvision; google2025aistudio), to adapt interfaces according to the user’s demonstrated behavior.
@@
To allow creative needs that step outside routine workflows, a scaffolded interface can act as a serendipity engine that encourages exploration. For instance, it can suggest unusual tool pairings based on deeper function relationships or highlight rarely used but relevant tools. This can be powered by LLMs, which use the user’s current state and project context to generate uncommon operations or workflows. We explore this potential by implementing scaffolded interface variations adapted to multiple creative demands, as shown in Appendix E.
@@
Scaffolded interface meta-layer for cross-software workflows. An enthusiasm by our expert users was using the scaffolded interface generation method in multiple software platforms. Experts highlighted the transferable task-aware, concept-driven UI design for task efficiency, using the scaffolded interface as a meta-layer to bridge different applications. Imagine that end users could design their own scaffolded interface built upon the tools they frequently use. Instead of learning a new interface for every software, they could keep this consistent interface that maps familiar tools to different underlying software. Such a interface could help improve cross-software workflows as many complex projects require switching between multiple applications, and create a unified experience, reducing the effort required by switching contexts. In addition, this scaffolded interface meta-layer can help users stay focused on learning fundamental concepts and operations instead of spending significant efforts on learning each software’s new interface. The implementation of this idea would require technical efforts to map concepts and operations to different underlying software, but the potential to simplify cross-platform workflows can make it interesting to explore for future research.

### v2

@@ 5.2. Study 2: Expert Users
For the study with experts (the advanced users who have developed their knowledge and skills through learning), we kept RQ1–RQ3 for a comparative analysis across beginner and expert users. In addition, two further research questions were formulated to explore our task-conditioned scaffolded UI’s impact on skill acquisition and our method’s generalizability.
@@
  • RQ4: How can the task-conditioned scaffolded UI potentially shape long-term skill development? This question invites expert perspectives on whether the task-conditioned scaffolded interface can help refine skills, adapt tools to new tasks, and strengthen concept understanding for solving new problems.
@@
  • RQ5: What is the potential impact of applying the task-conditioned scaffolded UI generation method to other software? This question asks for insights from experts familiar with additional professional creative software beyond Blender on the benefits and challenges of using task-conditioned scaffolded interface in different platforms.
@@
Experiment design. Each participant was assigned to one of two task conditions. Task 1 involved performing UV unwrapping on a 3D monkey head (Figure 8), an advanced texturing task, while Task 2 required constructing a human walk cycle that reflected personality, an open-ended animation task. We provided the 3D model for Task 1 and the rigged human character for Task 2 for experiment consistency. Task assignment was based on participant prior experience: only those with 3D model texturing experience were assigned to Task 1, and only experienced animators were assigned to Task 2. This allowed us to gather feedback grounded in the participants’ prior expertise to enable comparisons between their experience with the Blender interface and the task-conditioned scaffolded interface.
@@
Participant recruitment. We recruited eight experts (3 female, 5 male; mean age = 28.87, SD = 3.66) on Prolific, requiring at least one year of Blender experience and preferring additional proficiency in tools like Maya or 3ds Max. All other criteria for gender, age, language, and Prolific history matched those of Study 1. Participants (S2-P1 to S2-P8) were each assigned to either Task 1 or Task 2, and completed their task using the task-conditioned scaffolded interface, with four experts per task. Detailed demographics appear in Table 1.
Table 1. Study 2 (expert user study) participant demographic information.
This table summarizes the demographics of expert users, including their IDs, YoE, and experience in 3D modeling, animation, and additional software.
S2-P1 2 SketchUp
S2-P2 5 Maya, ZBrush, MotionBuilder
S2-P3 3 Maya
S2-P4 3 3ds Max, Mixamo, Unity, Unreal Engine
S2-P5 6 Unity, Maya
S2-P6 4 Tinkercad, 3ds Max, Maya
S2-P7 1 FreeCAD, Tinkercad
S2-P8 5 Substance 3D, After Effects
@@
Measurements. To evaluate RQs 1, 2, and 3, we used the same questionnaires as Study 1 and adapted the items to preference-based questions: each item was rated on a 7-point Likert scale (1 = strongly prefer our interface, 7 = strongly prefer the baseline interface). Additionally, we collected qualitative feedback by asking participants to elaborate on their experience with our interface regarding task performance, concept learning, and interface experience. For RQs 4 and 5, we gathered user qualitative data regarding our task-conditioned scaffolded UI’s impact on skill development and the potential of using the task-conditioned scaffolded UI in other software. Refer to Appendix B.4 for the open-ended questions for qualitative data collection.
@@
Procedure. We invited eligible participants to our study via Zoom. The pre-study session involving the consent and pre-study questionnaire had the same procedure as Study 1. Following this, we offered a brief introduction to the user’s task ($\sim$5 minutes). After participants verbally confirmed that they were clear about the task requirements, they used the task-conditioned scaffolded interface to work on the task for 15 minutes through remote desktop control. We recorded the desktop screen with participant consent and took notes on their actions. After the task, participants completed the post-study questionnaires and answered the open-ended questions ($\sim$10 minutes). The study concluded with a payment of 10 USD for completing the 30-minute study.
@@
Data analysis. The qualitative data was analyzed using thematic analysis (Braun and Clarke, 2006) with a deductive strategy. The authors reviewed the participant responses, coded the data to extract relevant information centered around the research questions, refined the codes, and categorized themes.
@@ 5.3. Study Design Remarks
Since our study with beginners aimed to explore their learning experience using our task-conditioned scaffolded UIs in a complex creative tool rather than making strong causal claims, not a large sample size was the main priority. We used a between-subjects design to avoid cross-condition contamination during learning and skill acquisition. In the study with experts, we prioritized qualitative depth over large-scale inference, and the eight users frequently interacted with Blender and were able to articulate pros and cons of our task-conditioned scaffolded interface clearly. Also, we observed signs of saturation during thematic analysis of the qualitative data.
@@ 6.1.1. Questionnaire Responses
Figure 9 shows the participant reported responses to the questionnaires. Asterisks in the figure indicate statistical significance derived from post-hoc pairwise comparisons between Ours (task-conditioned scaffolded interface) and Baseline (default Blender interface) for each task and measure. The complete two-way ANOVA results are reported in Appendix C. From the results, we see that our task-conditioned scaffolded interfaces have significantly improved task performance, supported user-perceived concept understanding and learning, and augmented user experience in using professional creative software compared to the default Blender interface.
@@
(c) User-perceived concept understanding and learning.
@@
Figure 9. Study 1 (beginner user study) participant responses using Ours (task-conditioned scaffolded interface) and Baseline (default Blender interface) for Tasks 1 and 2 (lower is better). Asterisks denote significant differences in post-hoc pairwise comparisons between interfaces within each measure (*: $p<.05$, **: $p<.01$, and ***: $p<.001$).
This figure shows the bar plots of Study 1 results. Our participants show clear preferences for our task-conditioned scaffolded interface across the measured four dimensions.
@@ 6.1.2. Task Completion Time
All participants successfully completed their tasks and generally spent less time using our task-conditioned scaffolded interface (Ours) compared to the default Blender interface (Baseline):
$\bar{t}_{Ours(Task1)}$ $=12.69\pm 1.49\textrm{ mins},$
$\bar{t}_{Baseline(Task1)}$ $=15.08\pm 2.34\textrm{ mins};$
$\bar{t}_{Ours(Task2)}$ $=20.39\pm 2.55\textrm{ mins},$
$\bar{t}_{Baseline(Task2)}$ $=24.95\pm 3.46\textrm{ mins}.$
Significant main effects of Interface ($F_{(1,30)}=14.78,p<.001,\eta^{2}_{p}=.33$) and Task ($F_{(1,30)}=94.71,p<.001,\eta^{2}_{p}=.76$) were observed. The interaction effect was not significant. Post-hoc analysis revealed no significant difference in completion time in Task 1, but users with our interface were significantly faster in Task 2 ($p<.05$), demonstrating the efficiency advantage of our task-conditioned scaffolded interface under higher task demands.
@@ 6.1.3. User Behavior
We observed that participants using our task-conditioned scaffolded interface (experiment group) progressively reduced their reliance on tutorial videos and increasingly navigated via the interface’s inherent workflow structure, echoed by participant self-reported improved workflow clarity, tool relevance, and reduced external help requirement. In contrast, participants using the default Blender interface (control group) consistently referred to the videos throughout the tasks. Consequently, the experiment group generally completed tasks faster than the control group. Additionally, when some tools were not mentioned in the tutorials but presented in our task-conditioned scaffolded interfaces, the experiment group explored them during or after their tasks, as echoed by participant self-reported tool and concept exploration. Conversely, the control group rarely explored beyond the tools covered in tutorials.
@@ 6.2.1. Questionnaire Responses
Figure 10 shows the participant reported preference scores for Ours (task-conditioned scaffolded interface) or Baseline (default Blender interface) on the measures of task performance, concept understanding and learning, and user interface experience. They had a clear preference for our interface across all dimensions regarding task performance. For concept understanding and learning, participants generally perceived our interface as more supportive for understanding concepts, relating tools to concepts and workflows, and motivating learning. However, they had mixed opinions on concept exploration (4 favored our interface, 1 was neutral, 3 favored the baseline). For the overall interface experience, consistent preference for our interface was observed except for tool exploration (5 participants preferred our interface, 3 favored the baseline).
Figure 10. Study 2 (expert user study) participant preferences for Ours (task-conditioned scaffolded interface) or Baseline (default Blender interface) regarding task performance, user-perceived concept understanding and learning, and user interface experience.
This figure shows the bar plots of Study 2 results. For most questions from the three evaluated aspects, the participants preferred our task-conditioned interface, except for ”Adjacent concept exploration” in perceived concept understanding and learning and ”Adjacent tool exploration” in user interface experience, the participants had mixed opinions.
@@ 6.2.2. Qualitative Results
Tradeoffs of supporting task performance and allowing workflow flexibility (RQ1). Participants generally appreciated task performance support by surfacing task-relevant tools (S2-P1, S2-P2, S2-P7, S2-P8). The task-conditioned scaffolded interface led to higher task efficiency and user focus, as S2-P2 said that, “I actually finished the UV task faster than expected.” However, participants also expressed concerns that the system’s assumptions sometimes misaligned with their preferred workflows. Some found the interface limiting when it failed to surface user-expected tools (S2-P4). Others perceived the guided structure forced a specific process, which may reduce workflow flexibility (S2-P6). These findings indicate the need for more user customization to support expert users who prefer flexible or iterative workflows.
@@
Gains of simplicity and focus vs. loss of familiarity (RQ3). Participants responded positively to the overall experience since our task-conditioned scaffolded interface was “clean, not overwhelming” (S2-P1), “task-centered” (S2-P4), “lightweight and approachable” (S2-P7), and with “limited distractions” (S2-P8). S2-P2 also noted that, “I can imagine a less fighting the UI if I were a newbie,” indicating reduced cognitive friction during software navigation for beginners. However, participants commented that the task-conditioned scaffolded interface may conflict with their familiarity with the default software layout. S2-P3 remarked that they “missed the old familiarity,” pointing to the tension when introducing novel interfaces.
Scaffolding’s impact on skill development (RQ4). Participants found that the task-conditioned scaffolded interface provided a supportive foundation for building practical skills, particularly for newcomers. For example, S2-P5 called it a “great stepping stone from beginner to intermediate,” and S2-P6 noted the interface made it “easier to build confidence in doing simple stuff.” These responses show that experts see our interface as helping beginners build confidence and gain skills while learning professional creative software.
Using task-conditioned scaffolded interfaces in broader contexts (RQ5). Participants with experience in other professional creative software recognized the underlying principles of task-conditioned scaffolded interface design as being transferable. For instance, S2-P8 stated that, “I think the knowledge about motion, timing, bone constraints translates to using other tools.” Building on such transferability, participants suggested potential applications and impacts of using task-conditioned scaffolded interfaces in other software. S2-P6 said that, “3ds Max can benefit from this. The modifier stack alone is complex enough. If tools were grouped by task and revealed based on what you’re actually doing, it’d cut down on confusion.” Participants also anticipated the benefits, e.g., “it could help new users get productive faster, and even experienced users might find it useful to stay focused” (S2-P6). These results suggest user enthusiasm for applying the task-conditioned scaffolded interface generation method more broadly to reduce interface complexity and enhance workflow efficiency.
@@ 6.2.3. User Behavior
We observed that expert users approached tasks flexibly, sometimes bypassing the step-by-step workflow stages in the task-conditioned scaffolded interface. For instance, S2-P6 switched between marking seams (stage 1) and unwrapping & editing (stage 2). They started with the “Select Edge Loops” tool in stage 1 to mark seams, used “Unwrap” in stage 2 to check the UV map, then went back to use “Edge Select” in stage 1 to refine the seams and repeated the process. This reflects their preference for multitasking and iterative workflows. Furthermore, when needed tools were not presented at their current complexity level, several participants (S2-P2, S2-P5, S2-P6, S2-P8) sometimes did not look for these tools by advancing complexity levels. Instead, they accessed tools via keyboard shortcuts or the default Blender interface—the operations they were familiar with. This aligns with their feedback that the task-conditioned scaffolded interface was less supportive for tool exploration compared to the experience of beginners.
@@ 7. Discussion
Based on the user study findings, here we discuss the opportunities in TaskLens, implications for design, TaskLens’s limitations, and future directions.
7.1. Opportunities in TaskLens
Task-aware adaptation to support software learners and educators. In our study, beginner users reported benefits in both task performance and user-perceived concept understanding with our task-conditioned scaffolded interface, showing its promise for instructional use. To better support learning for software learners, the interface can adapt its scaffolding strategy based on task complexity. For simple tasks, the interface can serve as a guided playground where future design could expand the available tools slightly beyond the current step to promote curiosity-driven exploration. For complex tasks, the focus shifts to practice and learning, so the interface can offer a smaller, essential toolset with clear links between domain concepts and software operations. It could also include best-practice tips or visual aids like video clips to help ensure key concepts and operations are grasped before moving on.
@@ 7.1. Opportunities in TaskLens
Scaffolded interface driven by productivity and creativity. Although our work focused on building task-conditioned scaffolded UIs to support software learning for beginners, our study with experts revealed that advanced users valued productivity through personalized workflows and creativity by flexible exploration. Therefore, scaffolded interfaces targeting at expert end users should evolve beyond structured guidance towards systems driven by these needs. To support productivity, a scaffolded interface could learn a user’s unique shortcuts, tool sequences, and niche commands, and optimize workflows by rearranging palettes for top tools or suggesting more efficient alternatives. Implementing this requires robust user activity logging, potentially through screen monitoring by in-application AI copilots (Cunningham, 2025; Google, 2025), to adapt interfaces according to the user’s demonstrated behavior.
@@
To allow creative needs that step outside routine workflows, a scaffolded interface can act as a serendipity engine that encourages exploration. For instance, it can suggest unusual tool pairings based on deeper function relationships or highlight rarely used but relevant tools. This can be powered by LLMs, which use the user’s current state and project context to generate uncommon operations or workflows. We explore this potential by implementing scaffolded interface variations adapted to multiple creative demands, as shown in Appendix D.
@@
Scaffolded interface meta-layer for cross-software workflows. An enthusiasm by our expert users was using the task-conditioned scaffolded interface generation method in multiple software platforms. Experts highlighted the transferable task-aware, concept-driven UI design for task efficiency, using the scaffolded interface as a meta-layer to bridge different applications. Imagine that end users could design their own scaffolded interface built upon the tools they frequently use. Instead of learning a new interface for every software, they could keep this consistent interface that maps familiar tools to different underlying software. Such an interface could help improve cross-software workflows as many complex projects require switching between multiple applications, and create a unified experience, reducing the effort required by switching contexts. In addition, this scaffolded interface meta-layer can help users stay focused on learning fundamental concepts and operations instead of spending significant efforts on learning each software’s new interface. We demonstrate this possibility by implementing scaffolded UIs for our user study tasks (UV unwrapping and walk cycle animation) in an additional professional software tool, Maya, and the results are available in Appendix E.
However, extending scaffolded UIs to broader domains and more varied tasks introduces foreseeable challenges. Tasks that are exploratory or open-ended in nature, e.g., freeform sculpting or experimental compositing, may resist the structured workflow decomposition that our current scaffolding relies on, and a constrained meta-layer could suppress the creative divergence these workflows demand. Additionally, maintaining semantic consistency across software platforms requires robust mappings between conceptually equivalent but operationally distinct tools, which may not always exist or may break as underlying applications evolve independently.
7.2. Implications for Design
Scaffolded UIs unify task instruction and interaction into a single design layer. A fundamental distinction between scaffolded UIs embedded within professional creative software and in-context assistance mechanisms, such as screen snippets highlighting relevant UI elements (Huang and Twidale, 2007) or procedural documentation guiding users through workflows (Bergman et al., 2005), lies in how each approach structures the relationship between task instruction and UI interaction. In-context assistance treats the application’s native interface and the instructional layer as separable concerns. Prior work has shown that with in-context assistance, users benefit most when visual guidance is tightly coupled with textual instructions, so that they can simultaneously understand what to do and where to do it within the interface (Huang and Twidale, 2007). Yet even when well-integrated, these overlays remain fundamentally external to the underlying UI, i.e., they annotate it rather than restructure it. Our scaffolded UI, by contrast, abstracts a focused, task-specific interface from the full application, unifying instruction and interaction within a single layer. Because the scaffolded UI is a malleable design surface, it affords the embedding of task guidance, learning materials, and other potential content directly into the interaction, rather than layering them on top of an interface whose complexity may itself be a barrier, especially for beginner users.
Tradeoffs between instructional integration and progressive skill transfer. This unification eliminates the cognitive overhead of context-switching between task instruction and execution. In the meantime, it introduces a tradeoff, where the scaffold UI could constrain the user’s exposure to the full application, while in-context assistance preserves the user’s engagement with the native interface. These tensions are echoed by our expert participants’ concerns about potential over-reliance on the scaffolded interface. Future design could include cues like UI minimaps or short animations showing where tools come from. The scaffolded interface could also fade as users gain skill, encouraging further transition to the native UI. Essentially, the design can adopt a transition from unified scaffolding to lightweight in-context cues, allowing the onboarding and learning benefits of scaffolding while ultimately fostering full interface fluency.